OpenStudy (anonymous):

a=3 b=-1 c=-10 use this to find the second solution of this equation -b-the square root of b^2-4ac all divided by 2a

pooja195 (pooja195):

The quadratic formula \[\huge~\rm~x=\frac{ -(-1) \pm \sqrt{(-1)^2-4(3)(-10)} }{ 2-(10) }\]

pooja195 (pooja195):

The quadratic formula \[\huge~\rm~x=\frac{ 1 \pm \sqrt{(-1)^2-4(3)(-10)} }{ 2-(10) }\]

pooja195 (pooja195):

Simplyfy

OpenStudy (anonymous):

a=3 not 10